OpenStudy (anonymous):

3 + radical 8 (2 - radical 5)

OpenStudy (turingtest):

\[3+\sqrt{8}(2-\sqrt{5})\] is the problem and you want it simplified?

OpenStudy (anonymous):

yeah. I don't understand how to get the simplified answer in a step by step manner.

OpenStudy (anonymous):

3+2 sqrt(2) (2-sqrt(5))

OpenStudy (turingtest):

as long as we're not missing parentheses anywhere, just distribute the sqrt(8):\[3+2\sqrt8-\sqrt{5\times8}=3+2\sqrt{4\times2}-\sqrt{4\times10}\]then simplify the radicals:\[=3+2(2\sqrt2)-2\sqrt10=3+4\sqrt2-2\sqrt10\]

OpenStudy (anonymous):

I don't understand the simplification process.

OpenStudy (turingtest):

ok say you have\[\sqrt40\]notice that this is the same as\[\sqrt{4\times10}\]since 4 is a perfect square we can pull it out of the radical:\[\sqrt{4\times10}=2\sqrt10\]like so. Want more examples/explanation?

OpenStudy (anonymous):

I understand that part. The part I didn't understand was after you put "then simplify the radicals"

OpenStudy (turingtest):

is it sqrt(8) or sqrt(2) in the firs part. you wrote it differently.

OpenStudy (turingtest):

3+2 sqrt(2) (2-sqrt(5)) 3 + radical 8 (2 - radical 5) are the two things you wrote, which is correct?

OpenStudy (anonymous):

3 + radical 8 (2 - radical 5)

OpenStudy (turingtest):

Well so after distributing the sqrt(8) we had\[3+2\sqrt8-\sqrt40=3+2(2\sqrt2)-2\sqrt10\]do you understand that step? It was just the same simplification you said you understood.

OpenStudy (anonymous):